Problem 3

A random sample Y1 , Y 2 , ⋯,Yn is drawn from a distribution whose probability density

function is given by: f (Y ) = βe− βY , Y  0 & β > 0

a). Obtain the maximum likelihood estimator (MLE) of β.

b). Given that Σ n

Y i = 25 ,

Σn

Yi2

= 50 , n = 50 calculate the maximum likelihood

i =1

i =1

estimate of β.

(3 point)

c). Using the same data as in part (b), test the null hypothesis that β =1against the alternative

hypothesis that β ≠1at 5% level of significance.



1
Expert's answer
2021-06-07T11:29:30-0400

(a) maximum likelihood estimator.


"\frac{df}{dy} = \beta""(\beta e""-\beta""Y)-""\beta"("\beta""e-\beta""Y)"



(b)

M= "\frac{1}{n}\sum" Xi (from 1to 50)

n= 50

Xi= 1

M="\frac{1}{50}= 0.02"


(c)T2"=\frac{1}{n}\sum"(Xi -M)2

"=0.02(1-0.02)= 0.0196"